OpenStudy (anonymous):

use the method of subsitution to solve the system of linear equations. 3x+y=-4 -2x+y=8

OpenStudy (anonymous):

convert the second equation to y=8+2x, replace in first so 3x+(8+2x)=-4

OpenStudy (anonymous):

solve for x , replace the found x value in any of the equations, and solve for y

OpenStudy (anonymous):

x=-12/5, y=-4+36/5

OpenStudy (anonymous):

how did you find the x and y value?

OpenStudy (anonymous):

the x value I found by solving 3x+(8+2x)=-4,>> 5x+8=-4, >> 5x=-12... the y i found by using the x value in the first equation